1/2

William Blake,  William Blake,
Der Tiger  The Tyger

tradukita de Walter A. Aue

Argiope-informo: La aspekto de tiu ĉi versio estas barita. Supozeble la aŭtoro decidis tion.

Tyger Tyger, burning bright, In the forests of the night; What immortal hand or eye, Could frame thy fearful symmetry?

In what distant deeps or skies Burnt the fire of thine eyes? On what wings dare he aspire? What the hand, dare sieze the fire?

And what shoulder and what art, Could twist the sinews of thy heart? And when thy heart began to beat, What dread hand? and what dread feet?

What the hammer? What the chain, In what furnace was thy brain? What the anvil? What dread grasp, Dare its deadly terrors clasp!

When the stars threw down their spears And water’d heaven with their tears: Did he smile his work to see? Did he who made the Lamb make thee?

…  …

2/2

Tyger Tyger, burning bright, In the forests of the night; What immortal hand or eye, Dare frame thy fearful symmetry?

Traduko de la Angla poemo “The Tyger ” de William Blake (∗1757-11-28 – †1827-08-12) en la Germanan de Walter A. Aue.

Arg-1048-2126 (2014-06-08 19:55:12)

Tiu ĉi poem-germanigo troviĝas en http: // myweb. dal. ca/ waue/ Trans/ Blake-Tiger. html .

Verkinto de tiu ĉi Angla poemo estas William Blake (∗1757-11-28 – †1827-08-12).

Arg-1048-2125 (2014-06-08 19:43:08)

Pri la poeto vidi la vikipediejon http: // eo. wikipedia. org/ wiki/ William_ Blake .

http://www.poezio.net/version?poem-id=1048&version-id=,2126,2125